OpenStudy (anonymous):

Solve the system, using substitution. Write the solution as an ordered pair. x = 2y - 2 and 3x - y = 14 A. (6, 4) B. (4, 6) C. (1, 0) D. (0, 1) E. no solution I put D is that correct? (:

OpenStudy (anonymous):

i got A here.

OpenStudy (anonymous):

Okay (: Are you sure this time? :D

OpenStudy (anonymous):

just substitute 2y-2 for the x in equation 3x-y=14 3(2y-2)-y=14 6y-6-y=14 5y-6=14 5y=20 y=4

OpenStudy (anonymous):

then plug in 4 in for y in the equation x = 2y - 2 x=2(4)-2 x=8-2 x=6

OpenStudy (anonymous):

Wait... wouldnt x = 4?

OpenStudy (anonymous):

ohh i didnt think about it that much ^^ thanks!!! :D

OpenStudy (anonymous):

lol. u r not thinking about these at all. u r in your own world. get into math world.

OpenStudy (anonymous):

I cant :( Every-time i do a problem i seem to get the wrong answer!! >.<

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

good luck. hope u start getting them not only getting the right answers. or Guessing multiple choice.